# Efficient Epistemic Uncertainty Estimation in Regression Ensemble Models Using Pairwise-Distance Estimators

### v3

@@ 7 Related Work
Researchers have employed Bayesian neural networks alongside information-based criteria for active learning in image classification problems (Gal et al., 2017; Kendall & Gal, 2017; Kirsch et al., 2019). In addition to these methods, others have proposed frameworks that utilize gradient embeddings in lieu of information-based criterion (Ash et al., 2019, 2021; Holzmüller et al., 2023). While the majority of these methods focus on image classification, a few methods have applications in regression (Ash et al., 2021; Holzmüller et al., 2023). In these cases, both Ash et al. (2021) and Holzmüller et al. (2023) only address 1D regression problems. In contrast, our research focuses on active learning in high-dimensional output space, up to 257 dimensions. In doing so, we address a gap in the literature which has importance in such fields as robotic locomotion, where gathering new data comes at high costs.
@@
In addition to Bayesian methods, ensembles have been harnessed for epistemic uncertainty estimation (Lakshminarayanan et al., 2017; Choi et al., 2018; Chua et al., 2018). Specifically related to our work, ensembles have been leveraged to quantify epistemic uncertainty in regression problems and active learning (Depeweg et al., 2018; Postels et al., 2020; Berry & Meger, 2023). Depeweg et al. (2018) employed Bayesian neural networks to model mixtures of Gaussians and demonstrated their ability to measure uncertainty in low-dimensional environments (1-2D). Building upon this foundation, Postels et al. (2020) and Berry & Meger (2023) extended the research by developing efficient Normalizing Flow (NF) ensemble models that effectively captured epistemic uncertainty. Our work advances this line of research by eliminating the need for sampling to estimate epistemic uncertainty, resulting in a faster and more effective method, especially in higher dimensions.
@@
Entropy estimators, which do not rely on sampling, is an active area of research (Jebara & Kondor, 2003; Jebara et al., 2004; Huber et al., 2008; Kolchinsky & Tracey, 2017). Kulak et al. (2021) and Kulak & Calinon (2021) demonstrated the utility of Pairwise-Distance Estimators (PaiDEs) within Bayesian contexts, employing PaiDEs to estimate conditional predictive posterior entropy. In contrast, our approach provides a more general estimate of epistemic uncertainty, as defined in Equation 1, which can be applied to both ensemble and Bayesian methods. Furthermore, our method is adaptable to flexible deep learning models, a capability that was previously unavailable in the approach presented by Kulak et al. (2021) and Kulak & Calinon (2021).
@@
Several methods have emerged in the literature for estimating epistemic uncertainty without relying on sampling techniques (Van Amersfoort et al., 2020; Charpentier et al., 2020). Both Van Amersfoort et al. (2020) and Charpentier et al. (2020) focus on classification tasks with 1D categorical outputs. Charpentier et al. (2021) extends the work of Charpentier et al. (2020) to regression tasks but is limited to modeling outputs as members of the exponential family. In contrast, our approach can handle more complex output distributions by directly considering the outputs from Normalizing Flows (NFs) and can be applied to a larger space of regression models. This flexibility is particularly valuable in scenarios involving intricate non-linear robotic dynamics, as demonstrated in our experiments.
@@ 8 Conclusions
In this study, we introduced two epistemic uncertainty estimators and applied them to expressive ensemble models. We depicted how our method can be used to more efficiently quantify uncertainty by leveraging closed-form pairwise-distance instead of sampling. This led to improvements in computational speed and accuracy, especially in larger dimensions. As learning becomes pervasive in high-dimensional and complex tasks in society, our method is well-placed to enable epistemic uncertainty awareness without undue computing overhead.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
The proofs follow from the steps of Kolchinsky & Tracey (2017).
@@
To show the upper bound from Equation 7 we use a derivation from Haussler & Opper (1997),
@@
where $E_{p_{i}}$ refers to the expectation when $Y$ is distributed as $p_{i}$ and $H(p_{i}\;\|\;p_{j})$ indicates the cross entropy. The inequality in line 2 follows from Hershey & Olsen (2007) and Paisley (2010). ∎
@@ Appendix B Compute and Hyper-parameter Details
The Nflows Base model employed one nonlinear transformation, $g$, with a single hidden layer containing 20 units, utilizing cubic spline flows as per (Durkan et al., 2019). The base network consisted of two hidden layers, each comprising 40 units with ReLU activation functions. It is important to note that all base distributions were Gaussian. The PNEs adopted an architecture of three hidden layers each with 50 units and ReLU activation functions. Model hyperparameters remained consistent across all experiments. Training was conducted using 16GB RAM on Intel Gold 6148 Skylake @ 2.4 GHz CPUs and NVidia V100SXM2 (16G memory) GPUs. For each experimental setting, PNEs and Nflows Base were executed with five ensemble components. The MC estimator sampled 1000 and 5000 points for Nflows Base and PNEs, respectively, for each $x$ conditioned on. The nflows library (Durkan et al., 2020) and the BMDAL library (Kirsch, 2023) were employed with minor modifications.
@@
where tr()($\cdot$)( ⋅ ) refers to the trace of a matrix.
@@ Appendix C Data
Regarding the multi-dimensional environments, namely Pendulum, Hopper, Ant, and Humanoid, the training sets and test sets were collected using different approaches. The training sets were obtained by applying a random policy, while the test sets were generated using an expert policy. This methodology was employed to ensure diversity between the training and test datasets. Notably, the OpenAI Gym library was utilized, with minor modifications (Brockman et al., 2016).
@@ Appendix D Additional Results
Figure 7: The bias introduced by PaiDEs for Nflows Base compared to the MC method on the 1D environments.
We provide a analysis of the bias introduced by PaiDEs in Figure 7. Note that in Figure 2 the epistemic uncertainty values mapped to 0-1 via a min-max normalization. Both plots show that PaiDEs introduce some bias in the 1D settings, though this is not a problem in an active learning setting as we only care about a point’s uncertainty relative to other points and this relationship is preserved. In addition to Table 1, we provide Figure 5 to show the entire active leaning curve and Table 2 detailing more acquisition batches. The pattern of PaiDEs outperforming or perform similarly to baselines holds across all environments and acquisition batches. An additional metric we evaluated on was log-likelihood, as the log-likelihood has been shown to be a proper scoring rule (Harakeh & Waslander, 2021). The results are shown in Figure 8, with 10 seeds being run and mean standard deviation being reported. As can be seen, the KL and Bhatt estimators perform similarly or better than the MC estimator on all environments. Note that due to the high-dimensional setting of Humanoid the log-likelihood for each estimator did not improve as data was added. This is due to the fact, that in order to calculate the log-likelihood of the ensemble one needs to calculate the likelihood of each individual model and then sum them together. In this process we are more likely to run into an issue of values being rounded to zero as we cannot store enough digits. As in Table 1, we have recorded certain acquisition batches in Table 3. Lastly, we provide a comparison on Hopper of the MC estimator to PaiDEs as the number samples drawn is increased in Figure 9.
@@
Env Acq. Batch Random BatchBALD BADGE BAIT LCMD Monte Carlo KL Bhatt
hetero 10 1.76 $\pm$ 0.17 6.81 $\pm$ 1.74 6.31 $\pm$ 2.05 6.89 $\pm$ 1.64 6.51 $\pm$ 1.22 1.48 $\pm$ 0.2 1.66 $\pm$ 0.2 1.56 $\pm$ 0.26
25 1.69 $\pm$ 0.45 6.98 $\pm$ 1.48 7.14 $\pm$ 0.72 7.64 $\pm$ 0.6 6.77 $\pm$ 0.75 1.44 $\pm$ 0.1 1.42 $\pm$ 0.12 1.42 $\pm$ 0.11
50 1.64 $\pm$ 0.22 7.26 $\pm$ 0.63 6.58 $\pm$ 0.91 7.46 $\pm$ 0.61 6.16 $\pm$ 0.88 1.45 $\pm$ 0.12 1.49 $\pm$ 0.29 1.42 $\pm$ 0.11
100 1.56 $\pm$ 0.14 6.62 $\pm$ 1.84 6.03 $\pm$ 1.68 7.39 $\pm$ 0.58 5.69 $\pm$ 1.03 1.54 $\pm$ 0.17 1.47 $\pm$ 0.15 1.55 $\pm$ 0.31
bimodal 10 8.4 $\pm$ 3.38 6.07 $\pm$ 0.13 6.07 $\pm$ 0.09 6.03 $\pm$ 0.05 6.19 $\pm$ 0.18 6.02 $\pm$ 0.05 6.01 $\pm$ 0.05 6.01 $\pm$ 0.05
25 6.1 $\pm$ 0.08 6.07 $\pm$ 0.09 6.17 $\pm$ 0.21 6.19 $\pm$ 0.23 6.14 $\pm$ 0.14 6.04 $\pm$ 0.04 6.02 $\pm$ 0.05 6.01 $\pm$ 0.04
50 6.57 $\pm$ 0.72 6.11 $\pm$ 0.09 6.14 $\pm$ 0.21 6.13 $\pm$ 0.12 6.17 $\pm$ 0.2 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.0 $\pm$ 0.04
100 6.4 $\pm$ 0.62 6.07 $\pm$ 0.12 6.25 $\pm$ 0.33 6.2 $\pm$ 0.41 6.24 $\pm$ 0.34 6.01 $\pm$ 0.04 6.01 $\pm$ 0.04 6.0 $\pm$ 0.04
Pendulum 10 0.28 $\pm$ 0.06 0.39 $\pm$ 0.24 0.64 $\pm$ 0.48 0.34 $\pm$ 0.19 0.34 $\pm$ 0.25 0.12 $\pm$ 0.04 0.15 $\pm$ 0.06 0.17 $\pm$ 0.08
25 0.22 $\pm$ 0.08 0.23 $\pm$ 0.3 0.15 $\pm$ 0.12 0.12 $\pm$ 0.09 0.11 $\pm$ 0.07 0.09 $\pm$ 0.03 0.09 $\pm$ 0.03 0.09 $\pm$ 0.02
50 0.18 $\pm$ 0.08 0.13 $\pm$ 0.1 0.14 $\pm$ 0.14 0.15 $\pm$ 0.12 0.11 $\pm$ 0.1 0.06 $\pm$ 0.03 0.06 $\pm$ 0.05 0.05 $\pm$ 0.01
100 0.15 $\pm$ 0.04 0.06 $\pm$ 0.03 0.07 $\pm$ 0.04 0.08 $\pm$ 0.05 0.05 $\pm$ 0.03 0.06 $\pm$ 0.01 0.05 $\pm$ 0.04 0.05 $\pm$ 0.01
Hopper 10 1.6 $\pm$ 0.19 1.03 $\pm$ 0.29 1.23 $\pm$ 0.44 1.05 $\pm$ 0.38 1.06 $\pm$ 0.32 0.55 $\pm$ 0.09 0.66 $\pm$ 0.08 0.69 $\pm$ 0.1
25 1.24 $\pm$ 0.26 0.64 $\pm$ 0.16 0.7 $\pm$ 0.21 0.63 $\pm$ 0.2 0.61 $\pm$ 0.12 0.36 $\pm$ 0.06 0.38 $\pm$ 0.05 0.39 $\pm$ 0.06
50 1.14 $\pm$ 0.16 0.53 $\pm$ 0.13 0.5 $\pm$ 0.1 0.53 $\pm$ 0.16 0.61 $\pm$ 0.26 0.31 $\pm$ 0.04 0.33 $\pm$ 0.03 0.34 $\pm$ 0.04
100 0.97 $\pm$ 0.2 0.62 $\pm$ 0.18 0.62 $\pm$ 0.21 0.68 $\pm$ 0.23 0.64 $\pm$ 0.2 0.32 $\pm$ 0.02 0.3 $\pm$ 0.03 0.29 $\pm$ 0.03
Ant 10 1.33 $\pm$ 0.1 2.23 $\pm$ 0.27 2.14 $\pm$ 0.29 2.21 $\pm$ 0.33 2.03 $\pm$ 0.22 1.24 $\pm$ 0.13 1.25 $\pm$ 0.07 1.29 $\pm$ 0.07
25 1.13 $\pm$ 0.09 2.21 $\pm$ 0.31 2.2 $\pm$ 0.32 2.13 $\pm$ 0.24 2.05 $\pm$ 0.19 1.13 $\pm$ 0.1 1.16 $\pm$ 0.07 1.2 $\pm$ 0.09
50 1.05 $\pm$ 0.05 2.21 $\pm$ 0.24 2.09 $\pm$ 0.42 1.93 $\pm$ 0.2 2.22 $\pm$ 0.58 1.05 $\pm$ 0.03 1.09 $\pm$ 0.04 1.1 $\pm$ 0.04
100 1.05 $\pm$ 0.1 1.93 $\pm$ 0.3 2.11 $\pm$ 0.26 1.91 $\pm$ 0.22 1.84 $\pm$ 0.17 1.01 $\pm$ 0.05 1.06 $\pm$ 0.08 1.1 $\pm$ 0.07
Humanoid 10 8.79 $\pm$ 1.17 11.51 $\pm$ 0.45 11.17 $\pm$ 0.2 10.97 $\pm$ 0.15 11.07 $\pm$ 0.26 10.69 $\pm$ 1.05 6.82 $\pm$ 1.34 7.5 $\pm$ 1.63
25 6.97 $\pm$ 1.48 11.4 $\pm$ 0.24 11.22 $\pm$ 0.37 11.01 $\pm$ 0.17 11.16 $\pm$ 0.2 10.65 $\pm$ 0.81 5.11 $\pm$ 1.34 5.3 $\pm$ 1.51
50 7.0 $\pm$ 1.34 11.3 $\pm$ 0.24 11.21 $\pm$ 0.31 11.07 $\pm$ 0.25 11.28 $\pm$ 0.19 10.94 $\pm$ 0.58 4.27 $\pm$ 0.94 4.08 $\pm$ 0.8
100 6.59 $\pm$ 1.54 11.28 $\pm$ 0.16 10.99 $\pm$ 0.33 11.01 $\pm$ 0.23 11.35 $\pm$ 0.16 10.71 $\pm$ 0.43 3.4 $\pm$ 0.48 3.37 $\pm$ 0.44

### v4

@@ 7 Related Work
Figure 5: The amount of time taken for Nflows Base for each estimator across the different settings (1, 3, 11, 27, 257 dimensions). Results are averaged over 10 seeds. MC results use $K=5000$ samples per acquisition candidate across all environments.
Figure 6: The amount of time taken for PaiDEs for Nflows Base as the number of ensemble components increases for Humanoid-v2. Results are averaged over 10 seeds.
Bayesian neural networks with information-based criteria are widely used for active learning in image classification [21, 32, 34], while others use gradient embeddings [3, 2]. Most focus on classification, with few methods extending to 1D regression [2]. Our work tackles active learning in high-dimensional regression, filling a gap relevant to robotic locomotion. Despite limitations of mutual information in classification [58], we demonstrate its value in high-dimensional regression.
@@
Ensembles have been harnessed for epistemic uncertainty estimation [39, 11, 12]. Specifically related to our work, ensembles have been leveraged to quantify epistemic uncertainty in regression problems and active learning [15, 48, 4]. Depeweg et al. [15] employed Bayesian neural networks to model mixtures of Gaussians and demonstrated their ability to measure uncertainty in low-dimensional environments (1-2D). Building upon this foundation, Postels et al. [48] and Berry and Meger [4] extended the research by developing efficient NF ensemble models that capture epistemic uncertainty. Our work advances this line of research by eliminating the need for sampling to estimate epistemic uncertainty, resulting in a faster and more effective method, especially in higher dimensions.
In addition to Bayesian neural networks and ensemble methods, the concept of Credal Bayesian Deep Learning has been introduced to enhance uncertainty quantification [7]. Furthermore, distributionally robust statistical verification has been proposed for high-dimensional autonomous systems using Imprecise Neural Networks, which provide uncertainty quantification guarantees [19].
@@
Entropy estimators, which do not rely on sampling, is an active area of research [30, 31, 28, 35]. Kulak et al. [37] and Kulak and Calinon [36] demonstrated the utility of PaiDEs within Bayesian contexts, employing PaiDEs to estimate conditional predictive posterior entropy. In contrast, our approach provides a more general estimate of epistemic uncertainty, as defined in Equation 1, which can be applied to both ensemble, Bayesian methods and deep learning models.
@@
Several methods have emerged in the literature for estimating epistemic uncertainty without relying on sampling techniques [56, 8]. Both Van Amersfoort et al. [56] and Charpentier et al. [8] focus on classification tasks with 1D categorical outputs. Charpentier et al. [9] extends the work of Charpentier et al. [8] to regression tasks but is limited to modeling outputs as members of the exponential family. In contrast, our approach can handle more complex output distributions by directly considering the outputs from NFs and can be applied to a larger space of regression models.
@@ 8 Conclusions
In this study, we introduced epistemic uncertainty estimators and applied them to Active Learning. We depicted how our method can be used to more efficiently quantify uncertainty by leveraging closed-form formulae instead of sampling. This led to improvements in computational speed and accuracy. As learning becomes pervasive in high-dimensional tasks in society, our method is well-placed to enable epistemic uncertainty awareness without unnecessary compute.
@@ Appendix A Proofs
The proofs follow from the steps of Kolchinsky and Tracey [35].
@@
To show the upper bound from Equation 6 we use a derivation from Haussler and Opper [23],
@@
where $E_{p_{i}}$ refers to the expectation when $Y$ is distributed as $p_{i}$ and $H(p_{i}\;\|\;p_{j})$ indicates the cross entropy. The inequality in line 2 follows from Hershey and Olsen [24] and Paisley [46]. ∎
A.1 PairEpEst-KL vs EPKL
In addition to the previous results from Kolchinsky and Tracey [35], we provide a unique proof to our paper demonstrating that the PairEpEst-KL estimator provides a strictly tighter upper bound on MI than the Expected Pairwise KL (EPKL) introduced by Malinin and Gales [42]. EPKL is another uncertainty measure that captures epistemic uncertainty by averaging the pairwise Kullback-Leibler divergences between predictive distributions of ensemble members. It quantifies how diverse the ensemble’s predictions are, reflecting the model’s uncertainty due to lack of knowledge or data. EPKL serves as an upper bound on MI. Our PairEpEst-KL estimator leverages a log-sum-exp formulation, yielding a sharper (tighter) upper bound on MI, which leads to more precise uncertainty quantification.
Specifically, consider an ensemble of $M$ predictive distributions $p_{1}(y),\ldots,p_{M}(y)$. The EPKL is defined as the average pairwise Kullback-Leibler divergence between distinct ensemble components:
$\mathrm{EPKL}:=\frac{1}{M(M-1)}\sum_{i\neq j}D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y)).$
Our PairEpEst-KL estimator, denoted by $B$, is given by
$B:=-\frac{1}{M}\sum_{i=1}^{M}\ln\left(\frac{1}{M}\sum_{j=1}^{M}\exp\big(-D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y))\big)\right).$
The following theorem formalizes the relationship between these two quantities and establishes that
$B<\mathrm{EPKL},$
meaning that PairEpEst-KL yields a strictly sharper upper bound on MI compared to EPKL.
Theorem A.1.
Let $p_{1}(y),\ldots,p_{M}(y)$ be probability distributions, and define
$A:=\frac{1}{M(M-1)}\sum_{i\neq j}D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y)),$
and
$B:=-\frac{1}{M}\sum_{i=1}^{M}\ln\left(\frac{1}{M}\sum_{j=1}^{M}\exp\big(-D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y))\big)\right).$
Then,
$B\leq\frac{1}{M^{2}}\sum_{i,j}D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y))<A,$
and in particular,
$B<A.$
Proof.
Define
$s_{ij}:=-D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y))\leq 0.$
By Jensen’s inequality applied to the concave function $\ln(\cdot)$, for each fixed $i$,
$\ln\left(\frac{1}{M}\sum_{j=1}^{M}e^{s_{ij}}\right)\geq\frac{1}{M}\sum_{j=1}^{M}s_{ij}.$
Multiplying both sides by 1-1 reverses the inequality:
$\ln\left(\frac{1}{M}\sum_{j=1}^{M}e^{s_{ij}}\right)\leq-\frac{1}{M}\sum_{j=1}^{M}s_{ij}.$
Substituting back for $s_{ij}$, we get
$\ln\left(\frac{1}{M}\sum_{j=1}^{M}\exp\big(-D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y))\big)\right)\leq\frac{1}{M}\sum_{j=1}^{M}D_{\mathrm{KL}}(p_{i}(y)\|p_{j}(y)).$
Averaging over $i$, it follows that
$B=\frac{1}{M}\sum_{i=1}^{M}\left[-\ln\left(\frac{1}{M}\sum_{j=1}^{M}e^{-D_{\mathrm{KL}}(p_{i}\|p_{j})}\right)\right]\leq\frac{1}{M^{2}}\sum_{i=1}^{M}\sum_{j=1}^{M}D_{\mathrm{KL}}(p_{i}\|p_{j}).$
Since $D_{\mathrm{KL}}(p_{i}\|p_{i})=0$, the double sum over all $i,j$ is equal to the sum over $i\neq j$:
$\sum_{i=1}^{M}\sum_{j=1}^{M}D_{\mathrm{KL}}(p_{i}\|p_{j})=\sum_{i\neq j}D_{\mathrm{KL}}(p_{i}\|p_{j}).$
Note that
$M^{2}>M(M-1)\implies\frac{1}{M^{2}}<\frac{1}{M(M-1)},$
so
$B\leq\frac{1}{M^{2}}\sum_{i\neq j}D_{\mathrm{KL}}(p_{i}\|p_{j})<\frac{1}{M(M-1)}\sum_{i\neq j}D_{\mathrm{KL}}(p_{i}\|p_{j})=A,$
which establishes the strict inequality
$B<A,$
completing the proof. ∎
@@ Appendix B Compute and Hyper-parameter Details
Algorithm 1 Active Learning Using PairEpEsts
Input: training dataset $\mathcal{D}_{train}$, oracle dataset $\mathcal{D}_{oracle}$, number of iterations $T$, sample size $S$, and batch size $B$
for i=1,2,…,$T$ do
  Randomly initialize each ensemble component $\theta_{j}$.
  Train our ensemble on $\mathcal{D}_{train}$ using bootstrapped samples.
  Sample $S$ points from $\mathcal{D}_{oracle}$ uniformly, $\mathcal{D}_{S}$.
  Calculate the top $B$ values according to Equation 9 from $\mathcal{D}_{S}$, $\mathcal{D}_{B}$.
  Add $\mathcal{D}_{B}$ to the training dataset while removing it from the oracle dataset, $\mathcal{D}_{train}=\mathcal{D}_{B}\cup\mathcal{D}_{train}$ and $\mathcal{D}_{oracle}=\mathcal{D}_{oracle}\setminus\mathcal{D}_{B}$
end for
 train final model on $\mathcal{D}_{train}$.
@@
The Nflows Base model employed one nonlinear transformation, $g$, with a single hidden layer containing 20 units, utilizing cubic spline flows as per [17]. The base network consisted of two hidden layers, each comprising 40 units with ReLU activation functions. It is important to note that all base distributions were Gaussian. The PNEs adopted an architecture of three hidden layers each with 50 units and ReLU activation functions. Model hyperparameters remained consistent across all experiments. The base distribution and the bijective mapping are trained simultaneously. Each base distribution network is mapped through the same bijective mapping, which is shared across all ensemble components. Every time a base distribution network takes a training step, the bijective mapping is updated accordingly. This approach ensures that the bijective mapping is consistent across all components of the ensemble. This methodology is similar to previous approaches, such as those proposed by Osband et al. [45], where shared mappings are utilized within ensemble-based models. Training was conducted using 16GB RAM on Intel Gold 6148 Skylake @ 2.4 GHz CPUs and NVidia V100SXM2 (16G memory) GPUs. For each experimental setting, PNEs and Nflows Base were executed with five ensemble components. The MC estimator sampled 1000 and 5000 points for Nflows Base and PNEs, respectively, for each $x$ conditioned on. The nflows library [18] was employed with minor modifications. Our code can be found at https://github.com/nwaftp23/pairflow-uncertainty.
@@
where tr()($\cdot$) refers to the trace of a matrix.
To complement the per-step acquisition runtimes reported in Figure 3, we also provide end-to-end active learning runtimes in Table 3 and Table 4. These results reflect the total wall-clock time required for a full active learning loop in each environment. These tables reinforce the efficiency advantage of PairEpEsts: across both architectures, KL and Bhatt estimators consistently require less total runtime than MC, particularly in higher-dimensional environments such as Ant and Humanoid. Notably, the runtime gap widens with dimensionality, reflecting the scalability of PairEpEsts relative to sampling-based methods.
To justify our choice of five ensemble components, we conducted an additional experiment analyzing the effect of ensemble size on active learning performance. This experiment was performed on the Hopper environment, with performance measured as RMSE at the final acquisition batch. Results are reported in Table 2. Performance improves markedly when increasing from 3 to 5 members, but remains stable for larger ensembles, suggesting that five components provide a strong balance between efficiency and accuracy.
Table 2: RMSE at the final acquisition batch for different ensemble sizes on Hopper.
3 4 5 7 10
KL 0.51 0.43 0.30 0.31 0.29
Bhatt 0.50 0.45 0.29 0.28 0.30
Table 3: Total active learning runtimes (minutes) for Nflows Base across environments.
Hetero Bimodal Pendulum Hopper Ant Humanoid
KL 58.14 58.22 76.02 77.06 80.21 96.28
Bhatt 57.72 58.04 76.86 76.94 78.54 95.58
MC 61.87 61.82 94.44 98.66 117.40 155.34
Table 4: Total active learning runtimes (minutes) for PNEs across environments.
Hetero Bimodal Pendulum Hopper Ant Humanoid
KL 22.68 22.50 28.44 28.60 30.58 35.24
Bhatt 22.50 22.36 28.24 28.88 30.14 36.04
MC 24.42 24.62 35.22 46.42 66.96 84.30
@@ Appendix C Data
Regarding the multi-dimensional environments, namely Pendulum, Hopper, Ant, and Humanoid, the training sets and test sets were collected using different approaches. The training sets were obtained by applying a random policy, while the test sets were generated using an expert policy. This methodology was employed to ensure diversity between the training and test datasets. Distribution shift is an inherent challenge in robotic dynamics, where data generated from different policies can vary significantly. These shifts, such as those observed in sim2real and imitation learning contexts. Notably, the OpenAI Gym library was utilized, with minor modifications [6].
@@ Appendix D Additional Results
Figure 8: The bias introduced by PairEpEsts for Nflows Base compared to the MC method on the 1D environments.
In addition to Table 1, we present Figure 3 showing the full active learning curves and Table 7 detailing additional acquisition batches. The trend of PairEpEsts outperforming or performing comparably to baselines is consistent across environments and acquisition batches. We also evaluated log-likelihood, a proper scoring rule [22], with results shown in Figure 9. Using 10 seeds, we report means and standard deviations. The PairEpEst-KL and PairEpEst-Bhatt estimators perform similarly or better than the MC estimator on all environments.
For the high-dimensional Humanoid setting, however, log-likelihood did not improve as data was added. This limitation arises because computing ensemble log-likelihood requires evaluating each component’s likelihood and summing them, a process prone to numerical underflow as values round to zero. As in Table 1, selected acquisition batches are reported in Table 8. We further compare the MC estimator and PairEpEsts on Hopper as the number of drawn samples increases (Figure 10). Additionally, we assessed the consistency of active learning rankings by computing Spearman’s rank correlation between the MC estimator and PairEpEsts. As shown in Table 5, correlations are close to 1 across all models and datasets, indicating strong preservation of relative ordering despite small biases. All correlations are statistically significant with extremely low p-values (maximum $p=1.11\times 10^{-83}$).
As summarized in Table 6, Monte Carlo methods achieve comparable runtime to PairEpEsts with roughly 10 samples in Hopper, and with even fewer samples in higher-dimensional environments such as Ant and Humanoid. This suggests that small-sample MC can, in principle, be competitive in terms of computational cost. However, PairEpEsts provide robust and consistent uncertainty estimates without the need for sample tuning, making them particularly advantageous in high-dimensional tasks where both runtime efficiency and estimator stability are critical. Moreover, the accuracy of MC estimates deteriorates as the number of samples decreases, whereas PairEpEsts maintain reliability at low computational cost.
Table 5: Spearman’s rank correlation coefficients comparing active learning rankings between MC and PairEpEsts, demonstrating that our estimators preserve relative ranking of points.
Model Dataset KL Corr. Bhatt Corr.
NFlows Base Bimodal 0.9958 0.9958
NFlows Base Hetero 0.9976 0.9986
PNEs Bimodal 0.9893 0.9893
PNEs Hetero 0.9943 0.9972
Table 6: Number of MC samples that matched the runtime of PairEpEsts across environments.
Environment Dimensionality MC Samples (Runtime Matched)
Hetero 1 500
Bimodal 1 500
Pendulum 3 100
Hopper 11 10
Ant 27 5
Humanoid 257 5
@@
Random BatchBALD BADGE BAIT MC (BALD) KL (ours) Bhatt (ours)
Env Acquisition Batch
[…]
bimodal 10 8.4 $\pm$ 3.38 7.37 $\pm$ 1.38 6.11 $\pm$ 0.1 6.91 $\pm$ 1.45 6.02 $\pm$ $\pm 0.05 6.01$$\pm$ 0.05
[…]
50 0.18 $\pm$ 0.08 0.15 $\pm$ 0.06 0.28 $\pm$ 0.14 0.17 $\pm$ 0.06 0.06 $\pm$ 0.03 0.06 $\pm$ $05±\pm 0.0$1
[…]
Hopper 10 1.6 $\pm$ 0.19 1.3 $\pm$ 0.26 1.45 $\pm$ 0.28 1.36 $\pm$ 0.19 $5±\pm 0$.09 0.66 $\pm$ 0.08 0.69 $\pm$ 0.1
[…]